# Dynamic and Distributed Routing in IoT Networks based on Multi-Objective Q-Learning

### v1

@@ Abstract
The last few decades have witnessed a rapid increase in IoT devices owing to their wide range of applications, such as smart healthcare monitoring systems, smart cities, and environmental monitoring. A critical task in IoT networks is sensing and transmitting information over the network. The IoT nodes gather data by sensing the environment and then transmit this data to a destination node via multi-hop communication, following some routing protocols. These protocols are usually designed to optimize possibly contradictory objectives, such as maximizing packet delivery ratio and energy efficiency. While most literature has focused on optimizing a static objective that remains unchanged, many real-world IoT applications require adapting to rapidly shifting priorities. For example, in monitoring systems, some transmissions are time-critical and require a high priority on low latency, while other transmissions are less urgent and instead prioritize energy efficiency. To meet such dynamic demands, we propose novel dynamic and distributed routing based on multiobjective Q-learning that can adapt to changes in preferences in real-time. Our algorithm builds on ideas from both multi-objective optimization and Q-learning. We also propose a novel greedy interpolation policy scheme to take near-optimal decisions for unexpected preference changes. The proposed scheme can approximate and utilize the Pareto-efficient solutions for dynamic preferences, thus utilizing past knowledge to adapt to unpredictable preferences quickly during runtime. Simulation results show that the proposed scheme outperforms state-of-the-art algorithms for various exploration strategies, preference variation patterns, and important metrics like overall reward, energy efficiency, and packet delivery ratio.
Index Terms:
Dynamic and Distributed Routing, Multiobjective, Q-Learning, Energy-efficiency, Internet of Things
@@ I Introduction
In recent years, rapid growth has been witnessed in Machine-to-Machine (M2M) communications and IoT networks owing to their wide range of applications. A report by CISCO on IoT growth [1] made an estimated count of over 29.3 billion IoT devices as of 2023, up from 18.4 billion in 2018. In IoT-based networks, the intelligent IoT nodes continuously monitor the environment and send the data to some destination node(s) when events occur. The destination node could be a sink node gathering data, or a base station, or another IoT node that needs to process the data. Due to their limited power, these nodes do not send data directly to the destination node; rather, the packets are sent via various relay nodes following certain routing protocols. To facilitate smooth and efficient data communication to meet the increasing demands of future IoT networks, designing intelligent routing protocols is important.
When designing intelligent routing protocols for IoT networks, multiple objectives must be considered. Among important objectives are maximizing the PDR, minimizing energy consumption, minimizing delay, ensuring security, and maximizing throughput. Prioritizing one objective in routing may come at the cost of another objective. For example, the most energy-efficient path for delivering packets from a source to a destination node may not be the one promising maximum PDR, as it may involve more link failures, network congestion, or selective forwarding attacks. If the routing protocols select such paths to save energy, it might adversely affect other objectives, e.g, the PDR. When the objectives are conflicting, no one solution exists that co-optimizes all objectives. In such cases, reasonable trade-offs based on the user’s preferences can be found by using the tools of multi-objective optimization [2]. This is often done by scaling the objectives with weights reflecting user and network preferences. These weighted objectives can then be used to construct a Pareto front, representing the set of optimal solutions where no single objective can be improved without worsening another, thereby providing a comprehensive view of the trade-offs involved.
In many real-wold IoT applications the objective is not static over time, but can be changing. Consider, for example, monitoring networks where certain information is time-critical, while other data requires transmission but is less sensitive to latency. In such cases we might prioritize low-latency for the time-sensitive transmissions, placing significant weight on the latency objective. Conversely, for less time-sensitive transmissions, we could prioritize other objectives, such as energy efficiency or security. As another example, consider a network with fluctuating energy supply. At certain times, when the energy are constrained, or nodes are running out of battery, there might be a need to prioritize energy conservation in the network, even if we may have to give lesser preference to other objectives. When energy is not scarce then we might prioritize other objectives. Such networks necessitate algorithms that can dynamically adjust to shifting objectives in real-time, ensuring optimal performance across varying conditions.
Future IoT networks will necessitate a new generation of distributed algorithms that are resource-aware and can optimize multiple objectives dynamically, adapting to the changing network preferences in real-time. Reinforcement Learning (RL) offers a promising solution to these problems owing to its ability to make intelligent responses to the environment based on experience of interacting with the environment. However, existing research (discussed in Section II) does not address dynamic objectives. Even studies considering multiple objectives typically scale these objectives without allowing for changes over time. The key contribution of this paper is to provide novel distributed algorithms that can learn to deal with multiple dynamic objective preferences in real-time.
In particular, our main contributions are:
@@
    We propose a novel dynamic preference Q-Learning routing algorithm that learns optimal decisions for multiple preferences in real-time. The algorithm adapts online to changing preferences, allowing it to continuously adjust to new objectives and conditions.
@@
    The proposed algorithm can be implemented in a distributed manner, with each node learning a local Q-table based on its actions, communicating only with neighbors in the routing network, and requiring no central information during training.
@@
    Extensive simulations compare our dynamic preference Q-Learning algorithm to six baseline schemes. We illustrate its performance using different exploration strategies, showing significant improvements in adaptability and routing efficiency over the baseline algorithms.
The paper is organized as follows: Section II discusses related work, Section III outlines the system model and problem formulation, Section IV details the proposed algorithm, Section V illustrates how the algorithm can be implemented in fully distributed manner, Section VI presents our simulation results, and Section VII concludes.
@@ II Related Works
Over the last two decades, there has been an upsurge in research related to developing efficient routing protocols for wireless networks and IoT. Some works present enhancements in Dijkstra’s algorithm for routing [3, 4]. However, such routing protocols are often associated with high local computations, storage requirements, and time complexities. Thus, it is inefficient to utilize them for resource-contained IoT networks today since they have to deliver quality services to users. Also, some other heuristic-based methods like sleep scheduling [5], fuzzy logic [3], trust-based or cluster-based approaches such as MANET, and VANET [6, 7], are also unable to meet the multiobjective expectations from IoT of the present and future. With the evolution of Machine Learning (ML), many ML-based approaches have been presented as promising solutions to routing-related optimization problems. Some of these routing algorithms are built on ML paradigms such as clustering [6], and evolutionary algorithms [8].
@@
RL, among other data-driven ML approaches, has attracted special attention since it provides us with the facility to learn optimal policies while interacting with the environment. Thus, many RL-based routing protocols have been proposed [9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29]. A direct utilization of RL for IoT is to rely upon a central controller for route planning and control. This has been made possible by the recent development of software-defined networking (SDN). In SDN-enabled IoT networks, a central controller has complete knowledge of the network topology and takes most of the control decisions in the network. The SDN-based centralized protocols face several challenges. Since the SDN controller plays a pivotal role in such protocols, it is an attractive target for malicious users to carry out DDoS attacks or other security attacks [30]. Such routing protocols are also vulnerable to single-point failures. Another critical issue in IoT networks is maintaining individual IoT nodes’ data privacy. Sharing all the data with a central controller can be both cost-inefficient and also pose a risk to data privacy. Moreover, as the size of IoT networks is increasing rapidly, scalability is a major issue with these centralized protocols [23].
@@
With more intelligent organization, it is feasible and beneficial to employ distributed versions of RL algorithms for routing. As an example, in tabular Q-learning, SDN-based and other centralized approaches rely on a centralized controller to keep the knowledge of the entire network, store the entire Q-table, and do the whole route planning as the RL agent. However, the Q-table can be split into smaller parts, with each node having its local small Q-tables with information about and interaction with their immediate neighbors. In this way, each node acts as an RL agent capable of making its own decisions. Some existing research works utilize this possibility and thus propose distributed RL-based routing protocols [15, 14]. Abadi et al. [13] utilize a sleep scheduling mechanism for enhancing the energy efficiency of their RL-based routing protocol. Such sleep scheduling is difficult to achieve in a fully distributed setup. Thus, they utilize a semi-distributed approach. They divide the network into clusters, with each cluster head acting as the local controller for the cluster. In this work, however, we showcase how a fully distributed RL-based scheme can be effectively applied while ensuring high energy efficiency and other vital objectives.
@@
While the RL-based routing protocols discussed so far work well for the static objective preference on which the model is trained, they cannot deal with changing preferences and objectives. However, in IoT networks, the user and network preferences keep changing in runtime. Existing RL-based approaches cannot quickly adapt to changes in the environment, and it takes time to relearn new policies suitable for new preferences. For example, Natarajan et al. [31] introduce an approach for distributed MORL with dynamic preferences. However, it does not approximate the entire Pareto front; rather, it learns from a limited number of random policies experienced in runtime. Further, it is limited to average reward RL algorithms (ARL) like R-learning and H-learning and not to discounted Q-learning. These approaches have some drawbacks that make them unsuitable for routing in IoT. It has been established in a previous work [32] that Q-learning consistently outperforms R-learning, even when both are evaluated with the same undiscounted performance metric. Additionally, R-learning is highly sensitive to the chosen exploration strategy. Interestingly, its performance notably worsens when using Boltzmann exploration. H-learning has the drawback that it consumes several times more CPU usage to converge than the model-free methods like Q-learning [33], thus unsuitable for routing in IoT networks. To overcome all the above-mentioned challenges, we propose a novel multi-objective Q-Learning-based distributed routing scheme that can utilize the Pareto frontier learned to find efficient routing schemes swiftly adapting to the dynamically varying preferences in the network.
@@ III-A Network Model and Routing Algorithms
A key objective of routing is to find the routing policy that optimizes specific metrics, such as energy consumption or PDR. In this context, each edge $(i,j)\in\mathcal{E}$ can be associated with a quantity $E(i,j)$ indicating, e.g., the energy needed to transmit data from $(i,j)\in\mathcal{E}$. We can now define the energy reward for non-terminal states $s=(i,j)$ and action $a$ as [^2] The reward is the negative of the true energy, as the objective is maximization.
@@ IV-A Q-Learning
A key advantage of Q-learning, which we leverage in our approach, is that it is an off-policy algorithm [35]. This means that the learning process is independent of the policy being followed during sample collection. In other words, Q-learning can learn the optimal Q-values regardless of the strategy used to explore the environment. As long as every state-action pair is explored sufficiently often, Q-learning is guaranteed to converge to the optimal Q-table. This property is particularly useful in multi-objective scenarios, since we can learn optimal Q-tables for many preferences simultaneously, irrespective of the actual routing policy being followed during the learning phase.
@@ IV-B Dynamic Preference Q-Learning
We illustrate this processes in Algorithm 1. We start by initializing the algorithm with the grid $\mathcal{B}$ and by initializing the Q-table $Q_{\beta}$ for each $\beta\in\mathcal{B}$. A straightforward initialization is to set all the Q-tables to zero, i.e., set $Q_{\beta}(s,a)=0$ for all $(s,a)\in\mathcal{S}\times\mathcal{A}$. After that we run a sequence of $M$ episodes, indexed my $m=1,2,\ldots,M$. Each episode begins with receiving an initial state $S^{0}=(i,j)$. The initial state indicates the next routing task where $i$ is the source and $j$ is the destination. We also receive a preference $\beta_{m}$ indicating how to balance the objectives during the routing task of the episode, as described in Eq. (4).
@@ IV-C Exploration Strategies
10:$\pi_{\beta}^{\texttt{Greedy}}(s):=\underset{a\in\mathcal{A}}{\text{argmax}}~{}Q_{\beta}^{\texttt{Int}}(s,a)$
@@
Simultaneous Exploration and Exploitation is an adaptive approach where exploration and exploitation occur together throughout the learning process. Early on, an epsilon-greedy strategy is used, meaning the agent selects random actions with a high probability (epsilon) to encourage exploration, while also exploiting the best-known actions at a smaller rate. Over time, epsilon is gradually reduced, placing less emphasis on exploration and more on exploitation as the agent learns. This method provides a more flexible balance between exploration and exploitation, allowing the agent to adapt as it gains knowledge. However, it can be slower to converge compared to purely exploiting after a set exploration phase

### v2

@@ Abstract
IoT networks often face conflicting routing goals such as maximizing packet delivery, minimizing delay, and conserving limited battery energy. These priorities can also change dynamically: for example, an emergency alert requires high reliability, while routine monitoring prioritizes energy efficiency to prolong network lifetime. Existing works, including many deep reinforcement learning approaches, are typically centralized and assume static objectives, making them slow to adapt when preferences shift. We propose a dynamic and fully distributed multi-objective Q-learning routing algorithm that learns multiple per-preference Q-tables in parallel and introduces a novel greedy interpolation policy to act near-optimally for unseen preferences without retraining or central coordination. A theoretical analysis further shows that the optimal value function is Lipschitz-continuous in the preference parameter, ensuring that the proposed greedy interpolation policy yields provably near-optimal behavior. Simulations show that our approach adapts in real time to shifting priorities and achieves up to 80–90% lower energy consumption and more than 2–5× higher cumulative rewards and packet delivery compared to six baseline protocols. These results demonstrate significant gains in adaptability, delivery, and efficiency for dynamic IoT environments.
@@ I Introduction
The Internet of Things (IoT) has grown rapidly in recent years, connecting billions of devices across domains such as smart healthcare, environmental monitoring, and smart cities. A fundamental task in IoT networks is to sense and transmit data over multi-hop wireless links to destination nodes such as sinks or base stations. Routing in such networks is challenging because nodes are energy-constrained, links are lossy, and traffic requirements vary widely across applications.
Designing efficient routing protocols for IoT requires balancing multiple, often contradictory objectives. For instance, the most energy-efficient path may traverse unreliable links and lead to high packet loss, while the most reliable or lowest-latency path may drain batteries quickly. Moreover, the relative importance of these objectives is not fixed: as an example, during time-critical events, latency and reliability are prioritized, whereas under normal conditions, energy conservation becomes more important to extend network lifetime. Future IoT systems must therefore support dynamic preference adaptation, where routing policies can adjust to shifting priorities in real time [liu2017performance, Vaishnav2024].
Classical approaches to multiobjective optimization often rely on weighted-sum formulations, where objectives are scaled by fixed weights and combined into a single cost function. While this provides a Pareto front of trade-offs, it assumes that preferences remain static [miettinen1999nonlinear]. In real-world IoT systems, however, priorities may change unpredictably during operation, making static weighting schemes insufficient.
Traditional routing solutions therefore fall short in dynamic IoT settings. Algorithms such as Dijkstra-based methods are computationally expensive for resource-limited devices. Heuristic schemes (e.g., fuzzy logic, clustering, or trust-based protocols) cannot adapt efficiently to changing objectives. Reinforcement learning (RL) approaches [kim2025deep, sharma2024optrisql, wang2024emor, 10419073, 10012574, 10608117, chakraborty2023intelligent, prabhu2023multiple, 9819969, 9488736, 9685191, 8935210] have emerged as promising alternatives, as they can learn routing decisions from interaction with the environment. However, most existing RL-based works (including deep RL methods) either (i) assume static objectives and retrain policies when preferences change, or (ii) rely on centralized controllers (e.g., SDN-based designs), which face problems of scalability, privacy, and single points of failure [frikha2021reinforcement]. This highlights the need to develop novel RL-based approaches for dynamic large-scale IoT deployments.
I-A Motivation
Real-world IoT deployments require routing protocols that are dynamic, distributed, and resource-aware. Specifically:
  • Dynamic adaptation: Objectives such as packet delivery and energy efficiency change in importance depending on the network state and application needs. A protocol must adapt quickly without retraining from scratch.
  • Efficient knowledge reuse: Since retraining is costly and slow, the system must leverage past learning to adapt instantly to new preference settings.
  • Distributed operation: Each IoT node has limited energy and knowledge of the network. Protocols must work without global controllers or complete topology knowledge, using only local neighbor information.
These requirements motivate the design of a routing scheme that can simultaneously prepare for multiple objective trade-offs, act near-optimally for unseen dynamic preferences, and operate in a fully distributed manner.
I-B Contributions and Novelty
This paper proposes a Dynamic Preference Q-Learning (DPQ-Learning) routing protocol that addresses the above challenges.
The key contributions of this work are summarized as follows:
@@ I-B Contributions and Novelty
    Distributed Multi-Objective Q-Learning Framework: A fully distributed tabular Q-learning scheme that learns multiple per-preference Q-tables in parallel from a single behavior stream, enabling efficient reuse of experience under IoT constraints.
@@
    A Novel Greedy Interpolation Policy for Dynamic Preferences: A lightweight interpolation mechanism that allows seamless adaptation to sudden changes in preference weights without retraining or centralized control.
@@
    Theoretical Guarantee of Near-Optimality: A formal analysis establishing that the interpolated policy remains provably near-optimal across the entire range of preferences, providing theoretical support for the algorithm’s real-time adaptability and stability.
The paper is organized as follows: Section II discusses related work, Section III outlines the system model and problem formulation, Section IV details the proposed algorithm, Section V illustrates how the algorithm can be implemented in fully distributed manner, Section VII presents our simulation results, and Section VIII concludes.
@@ II Related Works
TABLE I: Comparative summary of RL-based routing algorithms.
Ref Reinforcement Learning (RL) Method Multiobjective Optimization (MO) Method Adaptation to dynamic preferences MO Pareto Approximation Theoretical guarantees for MO method Fully Distributed
[kim2025deep] DRL + Intelligent neighbor node selection - - - -
[sharma2024optrisql, chakraborty2023intelligent, farag2021congestion] Tabular QL Scalarised linear reward - - -
[wang2024emor] RL with Actor-Critic architecture RL-learned weights (energy vs. progress) Likely - - -
[10419073] A3C Based Multistrategy RL Ratio-based reward - - - -
[10012574] Graph Neural Network (GNN) structure in Deep-RL - - - - -
[prabhu2023multiple] Multi-policy Deep-RL Ratio-based rewards - - - -
[8935210] Model free DRL (Actor-Critic) Linear scalarized reward - - - -
[kaur2021energy] Deep-RL (DRL) Linear scalarization (LS) function - - -
Our work Dynamic Preference Q-Learning with Greedy Interpolation
@@
Over the last two decades, there has been an upsurge in research related to developing efficient routing protocols for wireless networks and IoT. Some works present enhancements in Dijkstra’s algorithm for routing [5158843, 7435250]. However, such routing protocols are often associated with high local computations, storage requirements, and time complexities. Thus, it is inefficient to utilize them for resource-contrained IoT networks today since they have to deliver quality services to users. Also, some other heuristic-based methods like fuzzy logic [5158843], trust-based or cluster-based approaches such as MANET and VANET are also unable to meet the multiobjective expectations from IoT of the present and future. With the evolution of Machine Learning (ML), many ML-based approaches have been presented as promising solutions to routing-related optimization problems. Some of these routing algorithms are built on ML paradigms such as clustering, and evolutionary algorithms [10516314].
@@
RL, among other data-driven ML approaches, has attracted special attention since it provides us with the facility to learn optimal policies while interacting with the environment. Thus, many RL-based routing protocols have been proposed [sharma2024optrisql, chakraborty2023intelligent, farag2021congestion, wang2024emor, prabhu2023multiple, kaur2021energy, kim2025deep, 10419073, 10608117, 8935210, abadi2022rlbeep, bouzid2020efficient, guo2019optimizing, 9819969, 9634122, 9488736, cong2021deep, ergun2022reinforcement, 9685191, 10012574, phan2022deepair]. A direct utilization of RL for IoT is to rely upon a central controller for route planning and control. This has been made possible by the recent development of software-defined networking (SDN). In SDN-enabled IoT networks, a central controller has complete knowledge of the network topology and takes most of the control decisions in the network. The SDN-based centralized protocols [8935210, 10419073, 9634122, phan2022deepair, 9488736, cong2021deep, 10012574, 9819969] face several challenges. Since the SDN controller plays a pivotal role in such protocols, it is an attractive target for malicious users to carry out DDoS attacks or other security attacks [javanmardi2023sdn]. Such routing protocols are also vulnerable to single-point failures. Another critical issue in IoT is maintaining individual IoT nodes’ data privacy. Sharing all the data with a central controller can be both cost-inefficient and also pose a risk to data privacy. Moreover, as the size of IoT networks is increasing rapidly, scalability is a major issue with these centralized protocols [cong2021deep].
@@
With a more intelligent organization, it is feasible and beneficial to employ distributed versions of RL algorithms for routing. As an example, in tabular Q-learning, SDN-based and other centralized approaches rely on a centralized controller to keep the knowledge of the entire network, store the entire Q-table, and do the whole route planning as the RL agent. However, the Q-table can be split into smaller parts, with each node having its local small Q-tables with information about and interaction with their immediate neighbors. In this way, each node acts as an RL agent capable of making its own decisions. Some existing research works utilize this possibility and thus propose distributed RL-based routing protocols [guo2019optimizing, bouzid2020efficient]. Abadi et al. [abadi2022rlbeep] utilize a sleep scheduling mechanism for enhancing the energy efficiency of their RL-based routing protocol. Such sleep scheduling is difficult to achieve in a fully distributed setup. Thus, they utilize a semi-distributed approach. They divide the network into clusters, with each cluster head acting as the local controller for the cluster. In this work, however, we showcase how a fully distributed RL-based scheme can be effectively applied while ensuring high energy efficiency and other vital objectives. Some other distributed RL-based routing protocols [sharma2024optrisql, chakraborty2023intelligent, farag2021congestion, kim2025deep] have also been proposed, but they lack in dynamically adapting to changing preferences, as described below (Comparisons are summarized in Table. I).
@@
While the RL-based routing protocols discussed so far work well for the static objective preference on which the model is trained, they cannot deal with changing preferences and objectives. However, in IoT networks, the user and network preferences keep changing in runtime. Existing RL-based approaches cannot quickly adapt to changes in the environment, and it takes time to relearn new policies suitable for new preferences. For example, Natarajan et al. [natarajan2005dynamic] introduce an approach for distributed MORL with dynamic preferences. However, it does not approximate the entire Pareto front; rather, it learns from a limited number of random policies experienced in runtime. Further, it is limited to average reward RL algorithms (ARL) like R-learning and H-learning and not to discounted Q-learning. These approaches have some drawbacks that make them unsuitable for routing in IoT. It has been established in a previous work [mahadevan1994discount] that Q-learning consistently outperforms R-learning, even when both are evaluated with the same undiscounted performance metric. Additionally, R-learning is highly sensitive to the chosen exploration strategy. Interestingly, its performance notably worsens when using Boltzmann exploration. H-learning has the drawback that it consumes several times more CPU usage to converge than the model-free methods like Q-learning [tadepalli1994h], thus unsuitable for routing in IoT networks. To overcome all the above-mentioned challenges, we propose a novel multi-objective Q-Learning-based distributed routing scheme that can utilize the Pareto frontier learned to find efficient routing schemes swiftly adapting to the dynamically varying preferences in the network.
@@ III-A Network Model and Routing Algorithms
A key objective of routing is to find the routing policy that optimizes specific metrics, such as energy consumption or Packet Delivery Ratio (PDR). In this context, each edge $(i,j)\in\mathcal{E}$ can be associated with a quantity $E(i,j)$ indicating, e.g., the energy needed to transmit data from $(i,j)\in\mathcal{E}$. We can now define the energy reward for non-terminal states $s=(i,j)$ and action $a$ as [^2] The reward is the negative of the true energy, as the objective is maximization.
@@ IV-A Q-Learning
A key advantage of Q-learning, which we leverage in our approach, is that it is an off-policy algorithm [sutton1999reinforcement]. This means that the learning process is independent of the policy being followed during sample collection. In other words, Q-learning can learn the optimal Q-values regardless of the strategy used to explore the environment. As long as every state-action pair is explored sufficiently often, Q-learning is guaranteed to converge to the optimal Q-table. This property is particularly useful in multi-objective scenarios, since we can learn optimal Q-tables for many preferences simultaneously, irrespective of the actual routing policy being followed during the learning phase.
@@ IV-B Dynamic Preference Q-Learning
We illustrate this processes in Algorithm 1. We start by initializing the algorithm with the grid $\mathcal{B}$ and by initializing the Q-table $Q_{\beta}$ for each $\beta\in\mathcal{B}$. A straightforward initialization is to set all the Q-tables to zero, i.e., set $Q_{\beta}(s,a)=0$ for all $(s,a)\in\mathcal{S}\times\mathcal{A}$. After that we run a sequence of $M$ episodes, indexed by $m=1,2,\ldots,M$. Each episode begins with receiving an initial state $S^{0}=(i,j)$. The initial state indicates the next routing task where $i$ is the source and $j$ is the destination. We also receive a preference $\beta_{m}$ indicating how to balance the objectives during the routing task of the episode, as described in Eq. (4).
@@ IV-C Exploration Strategies
[…]
@@
Simultaneous Exploration and Exploitation is an adaptive approach where exploration and exploitation occur together throughout the learning process. Early on, an epsilon-greedy strategy is used, meaning the agent selects random actions with a high probability (epsilon) to encourage exploration, while also exploiting the best-known actions at a smaller rate. Over time, epsilon is gradually reduced, placing less emphasis on exploration and more on exploitation as the agent learns. This method provides a more flexible balance between exploration and exploitation, allowing the agent to adapt as it gains knowledge. However, it can be slower to converge compared to purely exploiting after a set exploration phase.